OpenStudy (anonymous):

How to solve for x ? log 4^(x+3) + log 4^(2-x) = 1

OpenStudy (anonymous):

use the identities : log(A.B) =log(A) + log(B) , and log(10)=1

OpenStudy (gorv):

wat is the base of log????

OpenStudy (gorv):

\[\log_{?} 4^(x+3+2-x)\]

OpenStudy (anonymous):

\[\log_{4^{\left( x+3 \right)}} + \log_{4^{\left( 2-x \right)}} = 1\]

OpenStudy (gorv):

@r.gupta04 now check the Q mate

OpenStudy (gorv):

change the base

OpenStudy (anonymous):

the one i mentioned was right...i:e its log to the base 10 by default ...and ln(x) is to the base e

OpenStudy (gorv):

make base of both log common

OpenStudy (gorv):

@r.gupta04 yeah sure u were right ...but he asked the Q which he dont know how to type

OpenStudy (anonymous):

guys the base of log is 4

OpenStudy (anonymous):

@gorv @r.gupta04

OpenStudy (gorv):

then \[\log_{4} ((x+3)*(2-x))=1\]=1

OpenStudy (gorv):

((x+3)*(2-x))=4^1 2x+6-x^2-3x=4 -x^2-x+6=4 -x^2-x+2=0 x^2+x-2=0 x^2+2x-x-2=0 (x+2)(x-1)=0 x=-2 or 1

OpenStudy (***[isuru]***):

then..\[\log_{4} x + 3 + \log_{4} 2-x = 1\] \[\log_{4} [ (x+3)(2-x)] =1\] which means \[4^{1} = (x+ 3)(2-x)\] \[4 = 6 -x -x^{2}\] \[x^{2} + x -2 = 0\] \[(x + 2)(x-1)=0\] \[x = -2 \ or \ x =1\]

OpenStudy (gorv):

@r. gupta check again

OpenStudy (***[isuru]***):

got it ? :-) I think u should have no prob in understanding cause u've seem to got lot of answers bro lol

OpenStudy (anonymous):

How do you guys cancel the \[\log_{4} \] ?

OpenStudy (anonymous):

4^(i wrote the expression thinking its - in the question....for this question

OpenStudy (***[isuru]***):